A trough is 75cm long and its ends in the form of an isosceles triangle have an altitude of 20cm 

and a base of 30cm. Water is flowing into the trough at the rate of 100cm3/sec. Find the rate at 

which the water level is rising when the water is 10cm deep?


1
Expert's answer
2022-04-19T03:37:26-0400


The volume of the water when the water is "h" cm deep:

"V=\frac{b\cdot L\cdot h}{2}"

From the similarity of triangles we can write:

"\frac{B}{H}=\frac{b}{h}" "\implies" "b=\frac{Bh}{H}"

Now the volume of the water:

"V=\frac{B\cdot L\cdot h^2}{2H}"

Let's find "\frac {dV}{dt}" :

"\frac {dV}{dt}=\frac{B\cdot L\cdot h}{H}\frac{dh}{dt}"

The rate at which the water level is rising:

"\frac{dh}{dt}=\frac{H}{B\cdot L\cdot h}\frac {dV}{dt}=\frac{20}{30\cdot 75\cdot 10}\cdot 100\approx0.089" cm/sec